OpenStudy (anonymous):

int from 1 to 2 [xlnx]

myininaya (myininaya):

\[\int\limits_{}^{}xln(x)dx=\frac{x^2}{2}\ln(x)-\int\limits_{}^{}\frac{x^2}{2} \cdot \frac{1}{x} dx+C\] \[=\frac{x^2}{2}\ln(x)-\int\limits_{}^{}\frac{x}{2}dx=\frac{x^2}{2}\ln(x)-\frac{1}{2} \cdot \frac{x^2}{2}+C\] \[=\frac{x^2}{2}(\ln(x)-\frac{1}{2})+C\] Check: \[[\frac{x^2}{2}(\ln(x)-\frac{1}{2})+C]'=x(\ln(x)-\frac{1}{2})+(\frac{1}{x}-0)\frac{x^2}{2}+0\] \[=x \ln(x)-\frac{1}{2}x+\frac{x}{2}=x \ln(x)\] now we have \[\int\limits_{1}^{2}x \ln(x) dx=[\frac{x^2}{2}(\ln(x)-\frac{1}{2})]_1^{2}\]

myininaya (myininaya):

\[=\frac{2^2}{2}(\ln(2)-\frac{1}{2})-\frac{1^2}{2}(\ln(1)-\frac{1}{2})=2(\ln(2)-\frac{1}{2})-\frac{1}{2}(0-\frac{1}{2})\] \[=2 \ln(2)-1+\frac{1}{4}=2\ln(2)-\frac{4}{4}+\frac{1}{4}=2 \ln(2)-\frac{3}{4}\]

OpenStudy (anonymous):

Cool! That's exactly what I got too..I was kind of tired of the repetitive algebra so I wanted to thrown some calculus in lol :)

OpenStudy (anonymous):

throw*

myininaya (myininaya):

try this one: \[\int\limits_{}^{}e^{\sqrt{3s+9}}ds\]

OpenStudy (anonymous):

okay

myininaya (myininaya):

god my cat just shedded all over me

myininaya (myininaya):

its time for a brushing

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

This is what I got:\[\ \frac{2\sqrt{3s+9}}{3} \ e ^{\sqrt{3s+9}}+C\]

OpenStudy (anonymous):

I see integration by parts was necessary in this case

myininaya (myininaya):

\[u^2=3s+9=>2u du=3 ds\] \[\frac{2}{3}\int\limits_{}^{}u e^u du=\frac{2}{3}(e^u u-\int\limits_{}^{}e^u du)+C=\frac{2}{3}e^u u-\frac{2}{3}e^u+C\] \[=\frac{2}{3}e^{\sqrt{3s+9}}\sqrt{3s+9}-\frac{2}{3}e^{\sqrt{3s+9}}+C\]

myininaya (myininaya):

\[=\frac{2}{3}e^{\sqrt{3s+9}}(\sqrt{3s+9}-1)+C\]

OpenStudy (anonymous):

yeah I see it know

myininaya (myininaya):

i think polpak showed me a way without integration by parts i can't remember

OpenStudy (anonymous):

hmm..

OpenStudy (anonymous):

Okay an

OpenStudy (anonymous):

okay another one

myininaya (myininaya):

another one? ok let me think

myininaya (myininaya):

\[\int\limits_{}^{}\ln(x+x^2)dx\]

OpenStudy (anonymous):

ok gimme a few minutes i'm actually going to eat i'll be back

OpenStudy (anonymous):

lol never mind I take that back I'll be back later

OpenStudy (anonymous):

Alright sorry i took long: Here is my answer:\[x \ln(x+x ^{2}) + \ln(x+1)-2x+C\]